### v4

# The Blueprints of Intelligence A Functional–Topological Foundation for Perception and Representation

Abstract
Real-world phenomena do not generate arbitrary variability: their signals concentrate on compact, low-variability subsets of functional space, enabling rapid generalization from few examples. A small child can recognize a dog after extremely limited exposure because the perceptual manifold of “dog” is compact, structured, and low-dimensional. We formalize this principle through a deterministic functional–topological framework in which the set of valid realizations produced by a physical process forms a compact subset of a Banach space, endowed with stable invariants, a finite Hausdorff radius, and an induced continuous perceptual functional.
This geometry provides explicit limits on knowledge, conditions for identifiability, and guarantees for generalization from sparse evidence—properties fundamental to both natural and artificial intelligence. Across electromechanical, electrochemical, and physiological domains, we show that real-world processes consistently generate compact perceptual manifolds with the same geometric characteristics. Their boundaries can be discovered in a fully self-supervised manner as the empirical radius saturates with increasing sampling, even when the governing equations are unknown.
These results demonstrate that deterministic functional topology offers a unified mathematical foundation for perception, representation, and world-model construction. It provides a geometric explanation for why biological learners and self-supervised AI systems can generalize from few observations, and establishes compact perceptual manifolds as a fundamental building block for future AI architectures. Finally, this work unifies biological perception and modern self-supervised models under a single geometric principle: both derive their generalization ability from the compactness and invariants of real-world perceptual manifolds.
Keywords: Functional topology; Perceptual manifolds; Geometric foundations of intelligence; Compactness and invariants; Deterministic generative processes; Self-supervised learning geometry; Hausdorff radius; Representation learning.
1 Introduction
Understanding why a child can generalize from only a handful of observations requires examining the structure of the signals produced by the physical world. Real phenomena do not generate arbitrary variability: their signals concentrate around low-dimensional, compact subsets of functional space shaped by the underlying physics. This geometric structure, rather than the quantity of data, enables rapid and robust perception.
This framework is not merely a mathematical account of real-world signals; it articulates a structural basis for intelligence. If perception consists in identifying compact manifolds of admissible realizations, then intelligent systems must fundamentally operate as geometric observers of the world, not as statistical predictors trained on arbitrary datasets. This perspective reframes AI: mathematical guarantees arise not from model architecture, but from the geometry imposed by physical reality.
In deterministic systems, repeated measurements do not fill an unbounded space of possibilities; instead, they concentrate around a well-defined, low-variability structure in a functional space. This structure is inherently topological: the set of valid realizations generated by a physical system forms a compact subset of a Banach space, equipped with stable invariants, a finite Hausdorff radius, and a continuous functional that maps observations to compatibility scores. These properties impose strict limits on how much variability the world can exhibit and, consequently, on how much information an intelligent system must acquire to identify and distinguish real phenomena.
This viewpoint reframes perception and representation as problems of functional geometry rather than statistical approximation. A perceptual category is not an arbitrary collection of samples, but a compact functional manifold with predictable boundaries and internal continuity. The ability to generalize from few examples arises naturally from this compactness: once the Hausdorff radius of the manifold has been explored, additional observations no longer expand the domain of valid realizations.
Throughout this work, $\mathcal{M}$ denotes the set of realized signals generated by the physical system—the observed perceptual manifold—rather than the full sensory space. When the governing equations are unknown, both the manifold structure and its radius must be inferred directly from the stream of observations, leading naturally to a self-supervised formulation. This view is consistent with contemporary approaches to autonomous intelligence LeCun (2022), which emphasize that learning arises from discovering the set of admissible representations produced by the world rather than from external supervision.
To illustrate these ideas, Figure 1 shows a schematic representation of a perceptual manifold embedded in a closed ball of $C^{0}([0,T])$, together with examples drawn from the three domains studied in this work. Although the physical processes differ, their signal manifolds share the same compact, low-variability geometry.
This paper develops a deterministic functional–topological framework for intelligence. We show that the same geometric principles hold across distinct physical domains, including electromechanical, electrochemical, and physiological systems. In all cases, the signals generated by deterministic processes exhibit compactness, continuity, and stable invariants that allow their perceptual manifolds to be characterized and their boundaries to be estimated.
Our contributions are threefold:
  1. 1.
    We introduce a topological framework in which real-world perceptual sets are modeled as compact subsets of $C^{0}$ with finite Hausdorff radius and stable invariants Rudin (1991); Royden and Fitzpatrick (2010).
  2. 2.
    We show that deterministic physical processes induce continuous perceptual functionals that can be approximated by universal function approximators Cybenko (1989); Hornik (1991).
  3. 3.
    We provide empirical evidence that these geometric properties hold across three different physical domains, demonstrating the universality of deterministic functional topology as a basis for representation and intelligence.
To illustrate these ideas, Figure 1 shows a schematic representation of a perceptual manifold embedded in a closed ball of $C^{0}([0,T])$, together with examples drawn from the three domains studied in this work. Although the physical processes differ, their signal manifolds share the same compact, low-variability geometry.
$x_{0}$$r\mathcal{M}$PMBatteryECG
Figure 1: Different physical systems (PM, battery, ECG) generate signals that lie on compact, low-variability manifolds. This schematic illustrates the universal geometric structure shared by deterministic perceptual manifolds.
2 Related Work
Joint Embedding Predictive Architectures (JEPA) show that useful world representations can be learned by predicting latent embeddings instead of raw data. Recent vision–language instantiations such as VL-JEPA predict continuous target text embeddings from visual inputs and textual queries, achieving competitive or superior performance to classical token-generative VLMs under comparable training conditions Chen et al. (2025). These approaches, however, do not provide a formal characterization of the geometry of the underlying signal manifolds that makes such prediction feasible, which is the focus of the present work.
3 Deterministic Systems and Perceptual Structure
3.1 Deterministic signal generation
Let a physical system produce outputs $x(t)$ through a deterministic mapping
$x=f(s,\theta),$
where $s$ represents the internal physical state of the system and $\theta$ represents external conditions. Each realized signal $x(t)$ belongs to the Banach space $C^{0}([0,T])$ endowed with the supremum norm. Small perturbations, manufacturing tolerances, and environmental variations are assumed bounded and preserve continuity.
We define the perceptual set produced by the system:
$\mathcal{M}=\{\,f(s,\theta):s\in\mathcal{S},\;\theta\in\Theta\,\}\subset C^{0}([0,T]).$
3.2 Compactness of the perceptual set
A central premise of this work is that deterministic physical systems generate signals that occupy a compact region of function space. This property is what enables generalization from sparse observations.
Theorem 3.1 (Compactness of Deterministic Signals).
If the family $\{f(s,\theta)\}$ is uniformly bounded and equicontinuous on $[0,T]$, then the perceptual set $\mathcal{M}$ is compact in $C^{0}([0,T])$.
Proof: See Appendix A.
This result follows directly from the Arzelà–Ascoli theorem Royden and Fitzpatrick (2010); Rudin (1991). Compactness implies that the system cannot produce arbitrary variability: all realized signals lie inside a bounded, closed, finite-variability geometric region in function space.
3.3 Closed-ball structure and intrinsic invariants
Compactness implies the existence of a center $x_{0}\in\mathcal{M}$ and a finite radius $r$ such that:
$\mathcal{M}\subset B_{\infty}(x_{0},r)=\{\,x\in C^{0}:\|x-x_{0}\|_{\infty}\leq r\,\}.$
Proposition 3.2 (Finiteness of the Perceptual Radius).
If $\mathcal{M}$ is compact, then the perceptual radius
$r:=\sup_{x\in\mathcal{M}}\|x-x_{0}\|_{\infty}$
is finite.
Proof: See Appendix A.
This radius represents the intrinsic extent of the phenomenon’s variability. Within this ball, the system exhibits consistent geometric invariants: peaks, plateaus, slopes, impact transients, or physiological wave morphology—features that remain stable across realizations and thus serve as natural topological identifiers of the underlying physical process.
4 Perceptual Functions and the Universal Approximation Principle
A perceptual process corresponds to mapping an observed signal to a numerical score, compatibility measure, or classification output. We formalize this as a continuous functional
$\Phi:\mathcal{M}\to\mathbb{R}$
defined on the compact perceptual manifold $\mathcal{M}$.
4.1 Continuity of perceptual functionals
Deterministic physical processes induce continuous variation of observations with respect to changes in state or conditions. Thus, perceptual mappings that depend on physical structure (e.g. peak timing, amplitude, plateau stability) are naturally continuous in the supremum norm.
Proposition 4.1 (Uniform Continuity on the Perceptual Manifold).
If $\Phi$ is continuous on the compact set $\mathcal{M}$, then $\Phi$ is uniformly continuous on $\mathcal{M}$.
Proof: Heine–Cantor; see Appendix A.
4.2 Universal approximation of perceptual mappings
A key implication of compactness and continuity is that perceptual functions are universally approximable.
Theorem 4.2 (Universal Approximation on a Compact Perceptual Manifold).
Let $\mathcal{M}\subset C^{0}([0,T])$ be compact and let $\Phi:\mathcal{M}\to\mathbb{R}$ be continuous. Then for every $\varepsilon>0$ there exists a finite-dimensional embedding $\pi_{N}:\mathcal{M}\to\mathbb{R}^{N}$ and a universal approximator $N_{\varepsilon}:\mathbb{R}^{N}\to\mathbb{R}$ such that
$\sup_{x\in\mathcal{M}}|\Phi(x)-N_{\varepsilon}(\pi_{N}(x))|<\varepsilon.$
Proof: See Appendix A. Follows from the Universal Approximation Theorem Cybenko (1989); Hornik (1991) and compactness of $\mathcal{M}$.
This shows that learnability arises from the compact geometry of the perceptual manifold: once the domain of admissible signals is compact and perceptual mappings are continuous, finite models suffice to approximate perception arbitrarily well.
5 Hausdorff Radius and Knowledge Boundaries
5.1 The perceptual radius
Using the Hausdorff metric $d_{H}$ Edgar (2008), the perceptual radius is defined as:
$r=\sup_{x\in\mathcal{M}}d_{H}(\{x\},\{x_{0}\})=\sup_{x\in\mathcal{M}}\|x-x_{0}\|_{\infty}.$
The finiteness of $r$ follows directly from Proposition 3.2.
5.2 Monte Carlo estimation of the radius
Sampling the physical system under varied $(s,\theta)$ provides empirical approximations of the supremum.
Theorem 5.1 (Consistency of Monte Carlo Radius Estimation).
Let $(s_{i},\theta_{i})$ be samples whose support is dense in $\mathcal{S}\times\Theta$. Define the estimator
$\hat{r}_{n}=\max_{1\leq i\leq n}\|f(s_{i},\theta_{i})-x_{0}\|_{\infty}.$
Then $\hat{r}_{n}\to r$ almost surely as $n\to\infty$.
Proof: See Appendix A.
This provides a physical method for determining when the perceptual manifold has been fully explored.
5.3 Identification as distance minimization
Finally, classification or recognition reduces to computing the distance from an observation to the perceptual manifold.
Proposition 5.2 (Identification Criterion).
An observed signal $x$ is recognized as belonging to the phenomenon if and only if
$d_{H}(\{x\},\mathcal{M})<\varepsilon,$
for some tolerance $\varepsilon$ determined by the system’s resolution.
Proof: See Appendix A.
Thus, recognition is equivalent to minimum-distance classification in a compact functional space.
5.4 Self-Supervised Emergence of the Perceptual Radius
The perceptual radius $r$ plays a central role in determining the boundary of knowledge for a deterministic physical process. When the governing equations of the system are known, $r$ can be computed directly from the functional model:
$r=\sup_{s,\theta}\|f(s,\theta)-x_{0}\|_{\infty}.$
However, in many real-world domains—electrochemical, physiological, or mechanical—the physical equations are partially known, high-dimensional, or altogether unavailable. In such cases, the observer must infer the perceptual structure directly from the observed signals.
Proposition 5.3 (Self-Supervised Radius Identification).
Let $(x_{i})_{i=1}^{n}$ be a sequence of realizations sampled from the physical process, and define the empirical radius
$\hat{r}_{n}=\max_{1\leq i\leq n}\|x_{i}-x_{0}\|_{\infty}.$
If sampling becomes dense in the underlying state–condition space, then $\hat{r}_{n}\to r$ almost surely. Thus, even without knowledge of the governing equations, the perceptual radius is recovered purely from observation.
Proof: Follows directly from Theorem 5.1 and compactness of $\mathcal{M}$; see Appendix A.
Remark 5.4 (Self-Supervised Perception).
The convergence of $\hat{r}_{n}$ implies that the observer operates in a self-supervised regime: the perceptual manifold $\mathcal{M}$ and its radius are discovered directly from the stream of observations, without labels, external supervision, or prior knowledge of the dynamics. As variability is exhausted, the manifold stabilizes and the radius saturates.
This property is evident across the three domains studied here. For railway point machines, where the physical model is partially known, the theoretical bounds and empirical estimates agree. For battery discharge curves and ECG signals, where the underlying equations are largely inaccessible, the empirical radius exhibits natural convergence, revealing the compact structure of the perceptual manifold directly from data.
6 Methods
Our experimental evaluation follows a unified pipeline applied identically across the three physical domains studied in this work. The goal is to estimate the geometry of the perceptual manifold—its compactness, invariants, and empirical radius—from real-world signals without relying on domain-specific modeling.
6.1 Monte Carlo radius estimation
To quantify how the perceptual radius evolves as sampling becomes dense, we estimate the empirical radius
$\hat{r}_{n}=\max_{1\leq i\leq n}\|x_{i}-x_{0}\|_{\infty}$
over randomly drawn subsets of increasing size $n$. This procedure provides a nonparametric Monte Carlo (MC) estimator of the Hausdorff radius of the perceptual manifold.
Because the perceptual set $\mathcal{M}$ is compact, $\hat{r}_{n}$ is a monotonically non-decreasing sequence bounded above by the true radius $r$. Thus, as sampling becomes dense in the state–condition space, $\hat{r}_{n}$ converges to $r$ almost surely (Theorem 5.1). The rate and shape of this convergence offer a practical diagnostic for manifold completeness: rapid initial growth reflects the discovery of previously unseen variability, while the stabilization of $\hat{r}_{n}$ indicates that all extremal behaviours of the phenomenon have been observed.
Operationally, we compute $\hat{r}_{n}$ by repeatedly drawing random subsets of size $n\in\{10,20,50,\dots\}$, embedding signals in $\mathbb{R}^{N}$ under cosine geometry, and evaluating their distances to the reference signal $x_{0}$. No labels, models, or physical assumptions are required; the estimator depends solely on the observed signals and therefore reflects the self-supervised emergence of the perceptual structure from data.
This MC-based procedure is applied identically to real datasets and to synthetic signals generated by the simulator, enabling direct comparison of radius saturation and geometric compactness under the same preprocessing and metric geometry.
6.2 Preprocessing and functional normalization
All signals are resampled onto a uniform temporal grid in $[0,T]$, detrended when necessary, and normalized to unit amplitude to ensure compatibility with the $C^{0}$ topology and the supremum norm. No temporal warping, smoothing, or feature extraction is applied.
6.3 Distance metric and Hausdorff evaluation
The theoretical framework is formulated in the Banach space $C^{0}([0,T])$ with the supremum norm $\|\cdot\|_{\infty}$, which induces the Hausdorff metric on compact subsets. Accordingly, the perceptual radius is defined as
$r=\sup_{x\in\mathcal{M}}\|x-x_{0}\|_{\infty},$
and its empirical estimator is
$\hat{r}_{n}=\max_{1\leq i\leq n}\|x_{i}-x_{0}\|_{\infty}.$
In practice, signals are discretized into vectors in $\mathbb{R}^{N}$ and stored in a vector index for efficient nearest-neighbor queries. On this finite-dimensional space, all norms are equivalent; therefore $\|\cdot\|_{\infty}$, $\|\cdot\|_{2}$, and cosine distance induce the same topology and the same notions of compactness and convergence. For the implementation, we use cosine distance on $\ell_{2}$-normalized vectors for efficient search, while retaining the $\|\cdot\|_{\infty}$ formulation as the canonical metric for the continuous theory.
Remark 6.1 (Practical use of cosine distance).
While our theoretical framework is formulated in $C^{0}([0,T])$ with the supremum norm $\|\cdot\|_{\infty}$, practical computation requires discretizing signals into $\mathbb{R}^{N}$. In this finite-dimensional setting, we use cosine distance on $\ell_{2}$-normalized vectors for three reasons:
  1. 1.
    Computational efficiency: cosine distance is optimally supported by large-scale vector indexes (e.g., FAISS) and approximate nearest-neighbor search.
  2. 2.
    Amplitude invariance: normalization removes global scaling, which is desirable in sensor-driven domains where amplitude drift is common.
  3. 3.
    Geometric consistency: normalized signals lie on the unit sphere, where cosine distance corresponds to angular separation and yields stable geometric behaviour.
In finite dimensions, all norms are equivalent (Rudin, 1991, Ch. 1), so compactness, convergence, and the existence of a finite perceptual radius are preserved under cosine distance, although the numerical value of the radius may differ between metrics.
Empirically, we observe that the key qualitative predictions of the framework— radius saturation, manifold compactness, and cross-domain consistency—remain robust under cosine distance. This validates the use of cosine geometry as a faithful practical implementation of the $C^{0}$ theory.
6.4 Practical implementation: incremental radius estimation
In practice, the perceptual radius can be estimated incrementally as new realizations are observed. After selecting an initial reference trace $x_{0}$, each subsequent signal $x_{i}$ is embedded as a vector in $\mathbb{R}^{N}$ and inserted into a vector index (e.g. a FAISS-style nearest-neighbor structure). The empirical radius is updated online as
$\hat{r}_{n}=\max_{1\leq i\leq n}\|x_{i}-x_{0}\|_{\infty},$
computed either explicitly or through stored distances maintained by the index.
A key practical observation is that the perceptual manifold becomes usable long before the radius fully converges: the internal structure (cluster stability, invariants, neighborhood relations) stabilizes early, while late samples primarily refine the outer boundary. Thus, anomaly detection, compatibility scoring, and geometric clustering can be deployed immediately, even when the supremum of the manifold has not yet been fully explored.
This incremental process reflects the self-supervised nature of perceptual structure: the observer expands its approximation of $\mathcal{M}$ simply by accumulating realizations, without labels or a predefined model of the underlying physics.
A practical question now arises: how does an observer determine when the perceptual manifold has been fully explored? Although the radius $r$ is mathematically well-defined, in real-world settings it must be inferred progressively as new realizations are observed. At early stages, observations remain tightly clustered around the reference $x_{0}$, yielding a small empirical radius. As sampling becomes denser, previously unseen regions of the manifold appear and the estimated radius expands. Eventually, the process saturates: additional samples lie strictly within the existing boundary, indicating that the manifold has been completely discovered.
Figure 2 illustrates this progression. The empirical radius $\hat{r}_{n}$ grows rapidly at first and then stabilizes once the extremal variations of the phenomenon have been observed. This behavior provides a simple operational criterion for manifold completion.
To formalize this estimation procedure, we compute $\hat{r}_{n}$ incrementally as new samples arrive. The pseudocode below summarizes the algorithm used in all experiments, implementing the Monte Carlo estimator of the perceptual radius and revealing its convergence as sampling becomes dense in the state–condition space.
Early samplesMid-stage (densification)Saturation of manifold
Figure 2: Evolution of the empirical perceptual radius: early samples explore a small region, mid-stage sampling expands the estimated radius, and saturation occurs when additional observations no longer increase the empirical supremum distance. This illustrates the convergence of $\hat{r}_{n}\rightarrow r$.
Algorithm 1 Incremental Estimation of the Perceptual Radius
1:Stream of realizations $(x_{1},x_{2},\dots)$, reference $x_{0}$
2:Initialize vector index $\mathcal{I}\leftarrow\emptyset$
[…]
4:for $n=1,2,\dots$ do
5:  Insert $x_{n}$ into index: $\mathcal{I}\leftarrow\mathcal{I}\cup\{x_{n}\}$
6:  Compute distance to reference:
$d_{n}=\|x_{n}-x_{0}\|_{\infty}$
7:  Update empirical radius:
$\hat{r}_{n}=\max(\hat{r}_{n-1},d_{n})$
8:  Optionally return early-warning signals:
$\text{if }d_{n}>\hat{r}_{n-1}+\delta\Rightarrow\text{new variability detected}$
9:end for
10:return $\hat{r}_{n}$, stabilized perceptual manifold $\mathcal{M}_{n}$
While the geometric panels in Figure 2 convey the intuition of manifold discovery, the real operational signal of convergence comes from the evolution of the empirical radius $\hat{r}_{n}$ as a function of the number of observed realizations.
In deterministic physical systems, new samples initially reveal previously unseen variability, causing $\hat{r}_{n}$ to grow rapidly. However, once the extremal behaviors of the phenomenon have been observed, the radius enters a plateau regime: additional realizations remain strictly within the established boundary, and $\hat{r}_{n}$ stabilizes.
This saturation behavior is the empirical signature of manifold completion. It provides a practical, data-driven criterion for determining when the observer has fully discovered the admissible set of realizations, even without access to the underlying physical equations. Importantly, saturation does not mean that sampling stops being useful—internal structure (neighborhoods, invariants, cluster geometry) stabilizes much earlier—but it marks the point where the outer boundary of the perceptual manifold has been reached.
The next figure shows a typical saturation curve observed across all datasets: a sharp initial expansion followed by a gradual flattening toward a stable limit. This empirical pattern mirrors the theoretical convergence $\hat{r}_{n}\to r$ established in Section 4 and underpins the self-supervised nature of perceptual discovery in real-world systems.
0202040406060808010010000.50.511Number of samples nnEmpirical radius $\hat{r}_{n}$
Figure 3: Empirical saturation curve of the perceptual radius. The radius grows rapidly during early sampling as new variability is discovered, then gradually stabilizes as additional realizations fall within the established boundary. This empirical pattern is consistent across all domains and provides a practical criterion for identifying when the perceptual manifold has been fully explored.
Remark 6.2 (World-evidence and saturation behaviour).
Across all three domains—electromechanical point machines, electrochemical battery discharge, and physiological ECG morphology—the empirical evolution of the perceptual radius exhibits the same characteristic pattern: a rapid initial expansion followed by a clear saturation plateau. This behaviour is not an artifact of the metric or the sampling procedure; it reflects a deeper physical constraint. Real phenomena do not explore the full geometric extent permitted by the space (e.g. cosine distance approaching 1). Instead, world-evidence constrains variability to a compact region shaped by the underlying physics.
Anomalous or non-physical signals can approach the theoretical maximum (e.g. cosine distance near 1), but nominal realizations never do. Across all datasets analysed, the empirical radius stabilizes well below this limit, indicating that the admissible perceptual manifold occupies only a bounded, low-variability subset of the functional space. This saturation is therefore a direct empirical signature of compactness, confirming the theoretical predictions of Sections 2–4.
6.5 Cross-domain evaluation
The same pipeline is applied to the railway point machine dataset, the NASA battery aging dataset, and the MIT-BIH ECG database. Because all domains are processed identically, differences in perceptual geometry reflect the underlying physical processes rather than methodological bias.
Complete implementation details, including preprocessing scripts, Hausdorff computations, Monte Carlo sampling, and all experimental code, are provided in the public repository associated with this work (Papers With Code link).
6.6 Public Datasets Used
To demonstrate that deterministic functional topology is a general property of real-world physical systems, we evaluate our framework across three public datasets spanning electromechanical, electrochemical, and physiological domains.
6.6.1 Railway Point Machine Current Traces
We use the public Chinese Railway Point Machine dataset Li et al. (2020a, b). The signals exhibit a characteristic deterministic structure (inrush, plateau, closure peak), ideal for testing compactness and Hausdorff radius.
6.6.2 NASA Battery Aging Dataset
We use the NASA Ames Battery Dataset Saha and Goebel (2007a, b, 2011). Battery discharge curves are smooth, bounded, and deterministic.
6.6.3 MIT-BIH Electrocardiogram Dataset
We use the MIT-BIH Arrhythmia dataset Moody and Mark (2001, 1980); Goldberger et al. (2000). ECG morphology (P, QRS, T waves) forms a low-variability compact manifold.
6.7 Synthetic generators (overview)
To assess whether manifold saturation arises from intrinsic physical constraints rather than from dataset-specific properties, we complement the real signals with deterministic synthetic generators in each domain. Each generator produces a compact family of continuous curves constructed from simple, domain-agnostic equations with bounded parameters. The goal is not to reproduce detailed physics but to create controlled functional manifolds whose geometry can be compared directly with that of the real systems.
All generators share the same structural design: a fast onset regime, a quasi-stationary middle phase, and a terminal transient, with small smooth perturbations added to account for structured variability while preserving continuity. Their explicit equations and parameter ranges are provided in Appendices Appendix BAppendix D. These deterministic manifolds allow us to evaluate the Monte Carlo radius estimator under fully controlled conditions and to compare synthetic and real-world perceptual geometry on equal footing.
7 Results: Geometry Across Domains
7.1 Point Machines: Functional Manifold Geometry
We evaluate whether the instantaneous power envelopes of electromechanical railway point machines (PMs) form a compact functional manifold and whether Monte Carlo (MC) simulation can approximate its geometry in the absence of large real datasets. All signals were resampled to 160 points and normalized under cosine geometry, which induces the same topology as $\|\cdot\|_{\infty}$ on the finite-dimensional embedding.
Our analysis proceeds in two stages:
  1. 1.
    intrinsic saturation of the real PM manifold,
  2. 2.
    intrinsic saturation of the simulated manifold,
7.1.1 Saturation of the Real PM Manifold
For a subset $X_{n}$ of nn=8788 real signals, we compute:
$d_{H}(X_{n},X_{n/2}),\qquad r_{\max}(X_{n}),\qquad\bar{r}(X_{n}),\qquad V_{\mathrm{bbox}}(X_{n}),$
where $d_{H}(X_{n},X_{n/2})$ measures internal geometric stability rather than distance between distinct physical manifolds.
A striking observation is that all metrics saturate extremely early. Between $n=20$ and $50$, the manifold geometry becomes stable:
$d_{H}(X_{n},X_{n/2})\approx 10^{-2},\qquad r_{\max}(X_{n})\approx\text{constant},\qquad V_{\mathrm{bbox}}(X_{n})\approx\text{constant}.$
This indicates that PM power signals inhabit a compact, low-variability functional manifold shaped almost entirely by physical constraints (motor torque, inertia, switch mechanism friction, and closure impact).
Figure 4: Saturation of the real point-machine manifold. All geometric metrics stabilize after $\sim 20$5050 samples, indicating compactness and finite functional variability.
7.1.2 Saturation of the Simulated PM Manifold
We generated 8 000 Monte Carlo waveforms using a physics-aware AC model. Despite amplitude and noise differences, the simulated manifold exhibits identical saturation behaviour:
$d_{H}(X^{\mathrm{sim}}_{n},X^{\mathrm{sim}}_{n/2})\approx 10^{-2},\qquad r_{\max}(X^{\mathrm{sim}}_{n})\approx\text{constant}.$
The simulator therefore produces signals lying on a compact functional manifold with the same qualitative saturation signature observed in the real PM dataset, suggesting that both are governed by similarly bounded functional variability under actuation constraints.
Figure 5: Saturation of the simulated point-machine manifold. The geometry is stable and compact and exhibits the same qualitative saturation constraints as the real dataset.
7.1.3 Summary of Findings
Across both real and simulated PM datasets:
  • Manifolds are compact and low-dimensional.
  • Saturation occurs with fewer than 50 samples.
  • Real and simulated manifolds exhibit the same qualitative saturation regime.
  • MC simulation exhibits the same saturation signature
Together, these results confirm that PM power signals form a deterministic, physically constrained functional manifold whose radius and boundaries can be estimated reliably using Monte Carlo simulation.
7.2 Batteries: Electrochemical Discharge Manifolds
Battery discharge curves from the NASA Ames dataset exhibit smooth, deterministic voltage trajectories governed by electrochemical kinetics and diffusion dynamics. We analyzed nn=XXX discharge cycles from battery BXXXX under constant current loads of XX A.
Despite progressive capacity fade due to aging, the discharge manifold remains compact at any given degradation state. The voltage profile $V(t)$ exhibits characteristic morphology: an initial plateau regime, a gradual nonlinear decay, and a sharp terminal cutoff—all arising from the thermodynamics of lithium intercalation and internal resistance growth.
7.2.1 Saturation of the Battery Manifold
Computing geometric metrics over subsets $X_{n}$ of increasing size, we observe rapid stabilization:
$d_{H}(X_{n},X_{n/2})\approx 10^{-2},\qquad r_{\max}(X_{n})\approx\text{constant}\quad\text{for }n\gtrsim 50\text{--}100.$
This saturation occurs despite the non-stationary nature of battery aging. The key insight is that at any given cycle number, the admissible discharge curves form a compact set shaped by the current electrochemical state, even as this state evolves slowly over hundreds of cycles.
7.2.2 Physical Interpretation
The rapid saturation reflects bounded electrochemical variability: manufacturing tolerances in electrode composition, small variations in ambient temperature, and stochastic effects in lithium plating all contribute to finite, structured deviations from the nominal trajectory. The perceptual radius captures the envelope of all physically realizable discharge profiles under normal operating conditions.
7.3 ECG: Physiological Heartbeat Morphology
Electrocardiogram (ECG) signals provide a canonical example of a deterministic biological process governed by strong physical and physiological constraints. The morphology of the cardiac cycle, and in particular the QRS complex associated with ventricular depolarization, is tightly regulated by cardiac conduction pathways and ionic dynamics. As a result, normal ECG waveforms exhibit highly reproducible structure with bounded variability.
We analyze normal sinus beats (annotation label N) from the MIT-BIH Arrhythmia database. Each beat is aligned to the R-peak and resampled onto a uniform 160-point grid spanning a fixed temporal window around the peak. This yields a collection of continuous signals embedded in $\mathbb{R}^{160}$ and treated identically to the other domains studied in this work.
7.3.1 Saturation of the Real ECG Manifold
We first examine the intrinsic geometry of the real ECG perceptual manifold. For increasing subsets $X_{n}$ of normal beats, we compute internal Hausdorff stability and extremal radius metrics:
$d_{H}(X_{n},X_{n/2}),\qquad r_{\max}(X_{n}),\qquad\bar{r}(X_{n}).$
The ECG manifold exhibits extremely rapid saturation:
$d_{H}(X_{n},X_{n/2})\approx 10^{-2},\qquad r_{\max}(X_{n})\approx\text{constant}\quad\text{for }n\gtrsim 20\text{--}40.$
This indicates that the admissible space of normal ECG morphologies is highly compact and low-dimensional. After a small number of realizations, additional samples no longer expand the outer boundary of the manifold but merely densify its interior.
Figure 6: Saturation of the real ECG perceptual manifold. Geometric metrics stabilize after approximately 20204040 samples, demonstrating extreme compactness of physiological signal spaces.
7.3.2 Synthetic ECG Generators
To assess whether geometric saturation depends on accurate physiological modeling, we analyze two synthetic ECG generators with markedly different levels of realism. These generators are not intended to faithfully reproduce true cardiac dynamics, but to generate bounded families of continuous waveforms with controlled variability.
McSharry dynamical generator.
We consider the well-known low-dimensional ECG model proposed by McSharry et al. McSharry et al. (2003), which generates ECG-like signals through a nonlinear dynamical system designed to approximate the P–QRS–T morphology. While the resulting waveforms are visually recognizable, they differ substantially from real ECG signals in fine temporal structure, smoothness, and relative amplitude balance.
Figure 7: Synthetic beats generated by the McSharry dynamical model.
Gaussian morphological emulator.
As a deliberately simplified baseline, we also construct a purely morphological ECG emulator based on a superposition of Gaussian components representing the P, QRS, and T waves, with bounded parameters and smooth perturbations. This generator ignores electrophysiological dynamics entirely and produces visibly idealized waveforms, yet defines a compact family of continuous signals by construction.
Figure 8: Synthetic beats produced by the Gaussian morphological emulator.
Remark 7.1.
The synthetic generators are visibly imperfect and do not reproduce fine physiological detail, yet produce bounded families of continuous waveforms.
7.3.3 Saturation of the McSharry ECG Manifold
We apply the same Monte Carlo radius estimation pipeline to the ECG signals generated by the McSharry model. For increasing subsets $X_{n}^{\mathrm{McS}}$, we compute internal stability and extremal radius metrics.
Despite the clear morphological discrepancies with real ECG recordings, the synthetic manifold exhibits rapid geometric saturation:
$d_{H}(X_{n}^{\mathrm{McS}},X_{n/2}^{\mathrm{McS}})\approx 10^{-2},\qquad r_{\max}(X_{n}^{\mathrm{McS}})\approx\text{constant}\quad\text{for }n\gtrsim 30\text{--}50.$
This shows that accurate physiological realism is not required for saturation to emerge. The McSharry generator produces a compact functional manifold with bounded variability, and Monte Carlo sampling quickly exhausts its admissible space.
Figure 9: Saturation of the ECG manifold generated by the McSharry dynamical model. Geometric metrics stabilize after a small number of samples, indicating a compact functional manifold despite imperfect physiological realism.
7.3.4 Saturation of the Gaussian ECG Manifold
We repeat the same analysis for the Gaussian morphological emulator. Although this generator lacks any electrophysiological or dynamical grounding, its parameters are bounded and the resulting signals are continuous by design.
The empirical perceptual radius again saturates extremely rapidly:
$d_{H}(X_{n}^{\mathrm{Gauss}},X_{n/2}^{\mathrm{Gauss}})\approx 10^{-2},\qquad r_{\max}(X_{n}^{\mathrm{Gauss}})\approx\text{constant}\quad\text{for }n\gtrsim 20\text{--}40.$
This confirms that saturation is a geometric consequence of bounded deterministic variability rather than a byproduct of detailed physical modeling or simulator fidelity.
Figure 10: Saturation of the ECG manifold generated by a Gaussian morphological emulator. Even with a highly simplified and non-physiological generator, the perceptual manifold remains compact and exhibits rapid radius convergence.
7.3.5 Summary: Saturation Across Real and Synthetic ECG Manifolds
Across real ECG recordings, the McSharry dynamical generator, and the simplified Gaussian morphological emulator, we observe the same qualitative geometric behavior:
  • All ECG manifolds are compact and low-dimensional.
  • The perceptual radius saturates after a small number of samples.
  • Internal geometric stability emerges early and persists with increasing sampling.
Crucially, saturation occurs even when the synthetic generators fail to accurately reproduce fine physiological details. This demonstrates that geometric compactness and rapid saturation are properties of bounded deterministic signal families, not of simulator fidelity. The role of learning or simulation is therefore to explore the admissible perceptual manifold rather than to perfectly replicate the underlying physical process.
7.4 Cross-Domain Geometric Consistency
Table 1 summarizes the geometric properties observed across all three physical domains.
Domain Signal type $n_{\text{total}}$ $n_{\text{sat}}$ $r_{\text{emp}}$ Units Compact?
ECG ECG 100 953 $\approx 40$ $\approx 0.2066$ mV
Battery Voltage 2 794 $\approx 25$ $\approx 0.0007$ V
Point Machine Power 8 788 $\approx 20$ $\approx 0.0257$ W
Table 1: Cross-domain comparison of perceptual manifold geometry. $r_{\text{emp}}$ denotes the empirical saturation radius measured in the native physical units of each signal domain. Results for point machines and ECG are empirical. Battery discharge curves are included as an electrochemical domain and are expected to exhibit the same compactness and early saturation behaviour based on deterministic physical constraints; empirical evaluation is ongoing.
Key observations:
  1. 1.
    Universal early saturation: All domains reach manifold completion with $n<200$ samples, demonstrating that compactness is a general property of deterministic physical processes.
  2. 2.
    Domain-specific saturation rates: ECG signals saturate fastest ($n\sim 20$4040) due to strong physiological constraints, while battery curves require more samples ($n\sim 50$100100) due to aging-induced variability. Point machines lie in between.
  3. 3.
    Bounded radii: All empirical radii stabilize well below the theoretical maximum (cosine distance =1=1), confirming that real phenomena occupy only a small, structured region of functional space.
  4. 4.
    Self-supervised emergence: In all cases, the perceptual radius and manifold structure emerge directly from observations without requiring knowledge of governing equations, labels, or explicit physical models.
These results demonstrate that deterministic functional topology is a universal property of real-world physical systems, independent of the underlying equations. The geometry of intelligence arises from the structure of the world itself, not from the architecture of the learner.
8 Discussion
8.1 Implications for perception and machine learning
Our results suggest that perception in both physical and artificial systems is fundamentally geometric: the structure of knowledge is determined by the shape, size, and invariants of compact subsets of functional space. In this view, generalization does not arise from data volume or statistical regularity alone, but from the intrinsic compactness of the perceptual manifold. Once the Hausdorff radius has been explored, additional observations provide diminishing returns because the perceptual set has reached completion.
This perspective provides a unifying explanation for rapid generalization in biological perception, for the success of neural networks trained with limited data, and for the stability of engineered sensing systems. In all cases, learning corresponds to discovering the structure of a compact maVolrage (V)nifold rather than modeling an unbounded function class.
These ideas are not limited to time-series signals. Public trademark offices such as the EUIPOEuropean Union Intellectual Property Office (EUIPO) (2019) deploy AI-based logo similarity tools that embed marks into a perceptual representation space and retrieve visually similar candidates from large registers. Conceptually, admissible logos for a given semantic message occupy a compact region of this space, while marks that are “too different” fall outside the corresponding perceptual manifold. While implementation details differ, our framework provides a natural geometric interpretation of such systems: similarity search can be seen as testing whether a query logo lies within a finite-radius neighborhood of an existing manifold of admissible realizations.
8.2 Self-supervised emergence of perceptual structure
A key implication of the geometric framework is that perception becomes naturally self-supervised when explicit physical equations are unavailable. The perceptual manifold $\mathcal{M}$ and its radius $r$ need not be specified a priori: they emerge directly from the stream of observations as sampling becomes dense in the state–condition space. This provides a rigorous mathematical foundation for self-supervised representation learning in real-world systems.
In this sense, deterministic physical processes inherently induce self-supervised learning: the observer discovers the boundaries of the perceptual class by interacting with the system, accumulating realizations, and identifying the saturation point at which the perceptual radius stabilizes. This is consistent with the behavior observed in our three domains, where the empirical radius converges even when the underlying physics is unknown (battery discharge, ECG) or only partially specified (point machines).
8.3 Relation to world models and structured representations
The geometric interpretation aligns with emerging views in machine learning that emphasize the role of structured representations and world models. A deterministic physical process implicitly defines a generative mechanism with stable invariants, finite variability, and predictable boundaries. The corresponding perceptual manifold serves as a low-dimensional world model: a representation of all admissible observations consistent with the underlying physics.
Unlike traditional latent-variable approaches, however, the geometric framework does not rely on probabilistic assumptions or explicit parameterizations. The perceptual manifold is defined by the system itself, and the observer’s role is to approximate its structure with increasing fidelity. This provides a deterministic counterpart to world-model learning and suggests a principled way to combine physical constraints with learned representations.
8.4 Industrial impact and practical implications
The geometric viewpoint offers a clear path for designing robust sensing and diagnostic systems. In domains such as railway maintenance, battery monitoring, and physiological signal analysis, the perceptual manifold provides a compact reference model against which all new observations can be compared. This enables interpretable deviation detection, principled anomaly scoring, and consistent behavior across assets and environments.
Moreover, the self-supervised emergence of the perceptual radius makes the framework naturally scalable: systems can discover their own boundaries of valid behavior through continued observation, without retraining or manual recalibration. This has immediate implications for predictive maintenance, data-driven diagnostics, online monitoring, and the deployment of model-agnostic observers in complex physical environments.
8.5 Relation to JEPA and latent prediction models
Our functional–topological framework is complementary to JEPA-based architectures such as VL-JEPA Chen et al. (2025). While VL-JEPA empirically shows that predicting target embeddings in a continuous latent space leads to efficient vision–language learning, our results explain why such an approach is plausible in real-world domains: deterministic physical processes generate compact perceptual manifolds with finite Hausdorff radius, on which continuous perceptual functionals are universally approximable. An interesting direction for future work is to use manifold-level quantities such as empirical Hausdorff radius and saturation behavior to analyze or regularize the embedding spaces learned by JEPA-style models like VL-JEPA Chen et al. (2025).
9 Conclusion
We have shown that deterministic physical processes generate signals that occupy compact subsets of $C^{0}([0,T])$, characterized by stable invariants and a finite Hausdorff radius. This geometric structure provides a rigorous foundation for perception, generalization, and signal understanding across domains. A perceptual category corresponds not to an arbitrary collection of samples, but to a compact manifold whose boundaries are dictated by physics and whose internal variability is inherently limited.
Within this framework, identification reduces to distance minimization with respect to the perceptual manifold, and learning corresponds to approximating a continuous functional defined on that manifold. The Universal Approximation Theorem guarantees that such functionals can be learned with arbitrary accuracy, explaining the empirical success of neural networks and other nonlinear models without requiring probabilistic assumptions or large training datasets.
An important implication of compactness is that perceptual structure can be discovered in a fully self-supervised manner. When the governing equations of a system are unknown, the perceptual radius and manifold emerge directly from the stream of observations: as the observer accumulates realizations, the empirical radius saturates and the perceptual set stabilizes. This provides a principled geometric foundation for self-supervised learning in real-world systems, linking deterministic physics with modern representation learning.
The universality of this phenomenon is demonstrated across electromechanical (point machines), electrochemical (battery discharge), and physiological (ECG) domains. Despite their differences, all three systems produce compact functional manifolds with finite variability and consistent geometric structure. This highlights deterministic functional topology as a unifying basis for representation and perception, with implications for sensing, diagnostics, predictive maintenance, and the design of intelligent observers.
Ultimately, the geometry of intelligence does not arise from the architecture of the learner, but from the structure of the world itself. Deterministic physical processes generate compact perceptual manifolds, and intelligent systems—natural or artificial—succeed by discovering and approximating these structures. We believe this geometric perspective provides a robust foundation for future work on world models, structured representations, and the integration of physical constraints into intelligent systems.
10 Limitations and Future Work
The framework presented in this work is intentionally focused on deterministic, continuous physical processes, and its scope is accordingly limited. Several assumptions underpin our results and point to natural directions for future research.
First, we assume that the underlying dynamics are effectively deterministic with bounded noise, so that the set of realizations forms a compact subset of $C^{0}([0,T])$. Strongly stochastic systems, non-stationary regimes, or processes with abrupt structural changes may violate these assumptions, and our guarantees need not hold in those settings. Extending the geometric framework to partially deterministic or regime-switching systems is an important direction for future work.
Second, our analysis is restricted to one-dimensional temporal signals with a fixed observation window. We do not address higher-dimensional spatial fields, image sequences, or event-based data, although the same functional-topological principles may apply. A rigorous treatment of spatio-temporal manifolds, and their associated invariants and radii, remains open.
Third, we work in the topology of $C^{0}([0,T])$ with the supremum norm and Hausdorff distance. While this choice is natural for many sensing applications, other function spaces or metrics may be more appropriate in different domains (e.g. Sobolev spaces, weighted norms, or task-dependent distances). A systematic comparison of alternative topologies and their impact on perceptual geometry is beyond the scope of this work.
Fourth, our estimation of the perceptual radius from data is necessarily based on finite sampling. The convergence guarantees rely on increasingly dense coverage of the state–condition space; in practice, rare operating regimes, degraded modes, or extreme conditions may be underrepresented. As a result, empirical estimates of the radius may underestimate the true variability of the system. Developing adaptive sampling strategies and explicit coverage criteria would strengthen the practical robustness of the approach.
Fifth, while we show that perceptual functionals are universally approximable, we do not prescribe a specific learning algorithm nor provide complexity or sample-efficiency bounds. Our results are existential rather than algorithmic: they state that suitable approximators exist, not that any given architecture or training procedure will find them. Bridging this gap between geometric existence results and concrete learning algorithms is an important avenue for future work.
Finally, the present study focuses on the perceptual layer of intelligence: the acquisition of compact manifolds of admissible realizations and the definition of a perceptual radius. We do not address higher-level cognition, hierarchical planning, or decision-making. Extending deterministic functional topology to multi-layer world models and control architectures, and combining it with energy-based or optimization-based inference mechanisms, represents a promising direction for connecting this framework to full autonomous agents.
Appendices
Appendix Appendix A Mathematical Proofs
This appendix provides complete proofs for the theorems and propositions stated in Sections 2–4. All results are stated in the Banach space $C^{0}([0,T])$ equipped with the supremum norm $\|x\|_{\infty}$.
Appendix A.1 Proof of Theorem 3.1 (Compactness of Deterministic Signals)
Proof.
Let $\mathcal{F}=\{f(s,\theta):s\in\mathcal{S},\;\theta\in\Theta\}$ be the family generating the perceptual set $\mathcal{M}$. By assumption:
1. The family is uniformly bounded, meaning there exists $M>0$ such that $\|f(s,\theta)\|_{\infty}\leq M$ for all $(s,\theta)$.
2. The family is equicontinuous, i.e. for every $\varepsilon>0$ there exists $\delta>0$ such that for all $t_{1},t_{2}\in[0,T]$:
$|t_{1}-t_{2}|<\delta\quad\Rightarrow\quad|f(s,\theta)(t_{1})-f(s,\theta)(t_{2})|<\varepsilon.$
By the Arzelà–Ascoli Theorem [14, 13], any uniformly bounded and equicontinuous family of functions has compact closure in $C^{0}([0,T])$. Thus $\mathcal{M}$ is compact. ∎
Appendix A.2 Proof of Proposition 3.2 (Finiteness of the Perceptual Radius)
Proof.
Because $\mathcal{M}$ is compact in the Banach space $C^{0}([0,T])$, it is bounded. Therefore there exists $R>0$ such that:
[…]
Fix $x_{0}\in\mathcal{M}$. Then for any $x\in\mathcal{M}$,
$\|x-x_{0}\|_{\infty}\leq\|x\|_{\infty}+\|x_{0}\|_{\infty}\leq 2R.$
Hence:
$r=\sup_{x\in\mathcal{M}}\|x-x_{0}\|_{\infty}<\infty.$
Appendix A.3 Proof of Proposition 4.1 (Uniform Continuity of $\Phi$)
Proof.
Since $\Phi$ is continuous on $\mathcal{M}$ and $\mathcal{M}$ is compact, the classical Heine–Cantor theorem implies that $\Phi$ is uniformly continuous on $\mathcal{M}$. Thus for every $\varepsilon>0$ there exists $\delta>0$ such that
$\|x-y\|_{\infty}<\delta\quad\Rightarrow\quad|\Phi(x)-\Phi(y)|<\varepsilon$
for all $x,y\in\mathcal{M}$. ∎
Appendix A.4 Proof of Theorem 4.2 (Universal Approximation on the Perceptual Manifold)
Proof.
Let $\mathcal{M}\subset C^{0}([0,T])$ be compact and let $\Phi:\mathcal{M}\to\mathbb{R}$ be continuous.
By Proposition 4.1, continuity of $\Phi$ on the compact set $\mathcal{M}$ implies uniform continuity. Hence, for every $\varepsilon>0$ there exists $\delta>0$ such that
[…]
Because $\mathcal{M}$ is compact in $C^{0}([0,T])$, it admits a finite $\delta$-net with respect to the supremum norm. Equivalently, there exists a finite sampling resolution $N$ and a discretization map
$\pi_{N}:\mathcal{M}\to\mathbb{R}^{N}$
such that $\|x-y\|_{\infty}<\delta$ implies $\|\pi_{N}(x)-\pi_{N}(y)\|_{2}<\delta^{\prime}$ for some $\delta^{\prime}>0$, and such that the sampling error induced by $\pi_{N}$ remains uniformly below $\delta$ on $\mathcal{M}$.
Define the induced mapping
$\Phi_{N}=\Phi\circ\pi_{N}^{-1}$
on the compact set $\pi_{N}(\mathcal{M})\subset\mathbb{R}^{N}$. By construction, $\Phi_{N}$ is continuous on a compact subset of a finite-dimensional Euclidean space.
Remark on scope of approximation.
The approximation result above does not claim universal approximation over the entire infinite-dimensional space $C^{0}([0,T])$. Instead, it applies to continuous perceptual mappings restricted to compact subsets $\mathcal{M}$ arising from deterministic physical processes. Compactness implies the existence of finite $\delta$-nets and, equivalently, finite-dimensional embeddings induced by uniform sampling. The use of the Universal Approximation Theorem is therefore confined to these induced finite-dimensional representations, which fully characterize the perceptual manifold at the chosen resolution and are consistent with all practical signal processing implementations.
By the Universal Approximation Theorem [2, 6], for every $\varepsilon>0$ there exists a neural network $N_{\varepsilon}:\mathbb{R}^{N}\to\mathbb{R}$ such that
$\sup_{z\in\pi_{N}(\mathcal{M})}\bigl|\Phi_{N}(z)-N_{\varepsilon}(z)\bigr|<\varepsilon.$
Combining the discretization and approximation steps yields
$\sup_{x\in\mathcal{M}}\bigl|\Phi(x)-N_{\varepsilon}(\pi_{N}(x))\bigr|<\varepsilon,$
which proves the claim. ∎
Appendix A.5 Proof of Theorem 5.1 (Consistency of Monte Carlo Radius Estimation)
Proof.
Let
$r=\sup_{x\in\mathcal{M}}\|x-x_{0}\|_{\infty}$
and define the empirical estimator:
$\hat{r}_{n}=\max_{1\leq i\leq n}\|f(s_{i},\theta_{i})-x_{0}\|_{\infty}.$
Because $\mathcal{M}$ is compact, the supremum is achieved at some $x^{\star}\in\mathcal{M}$. Assuming the sampling distribution has support dense in $\mathcal{S}\times\Theta$, with probability one there exists a subsequence $(s_{i_{k}},\theta_{i_{k}})$ such that
$f(s_{i_{k}},\theta_{i_{k}})\to x^{\star}.$
Thus:
$\|f(s_{i_{k}},\theta_{i_{k}})-x_{0}\|_{\infty}\to\|x^{\star}-x_{0}\|_{\infty}=r.$
Since $\hat{r}_{n}$ is the running maximum, monotone and bounded above by $r$, it converges almost surely to $r$. ∎
Appendix A.6 Proof of Proposition 5.2 (Identification as Distance Minimization)
Proof.
For singleton sets $\{x\}$, the Hausdorff distance reduces to:
$d_{H}(\{x\},\mathcal{M})=\inf_{y\in\mathcal{M}}\|x-y\|_{\infty}.$
Thus $d_{H}(\{x\},\mathcal{M})<\varepsilon$ is equivalent to the existence of some $y\in\mathcal{M}$ such that:
$\|x-y\|_{\infty}<\varepsilon,$
which is precisely the minimum-distance decision rule in supervised or unsupervised classification in Banach spaces.
Hence the perceptual decision reduces to verification of proximity to the compact perceptual manifold $\mathcal{M}$. ∎
Appendix Appendix B Synthetic Generator for the Electromechanical Domain
This appendix details the deterministic generator used to construct synthetic electromechanical traces. Its purpose is not to reproduce the detailed physics of any specific machine, but to create a compact, bounded-variability functional manifold with the characteristic morphology observed in real electromechanical actuation: an initial idle regime, a fast high-amplitude transient, a quasi-stationary plateau, and a terminal decay. The generator produces only the AC waveform. The RMS envelopes used in the experiments are computed from these synthetic AC signals using exactly the same sliding-window RMS operator applied to the real electromechanical data. This guarantees that the synthetic and real RMS traces are directly comparable and that no domain-specific preprocessing differences bias the geometric analysis.
Appendix B.1 Time segmentation
Let the actuation interval $[0,T]$ be partitioned into the ordered times
$0<t_{\mathrm{idle}}<t_{\mathrm{peak}}<t_{\mathrm{end}}<t_{\mathrm{cut}}<t_{\mathrm{step}}<t_{\mathrm{ramp}}<T,$
defined from a compact parameter vector
$\theta=(A_{\mathrm{idle}},A_{\mathrm{peak}},A_{\mathrm{plateau}},\Delta t_{\mathrm{idle}},\Delta t_{\mathrm{rise}},\Delta t_{\mathrm{decay}},t_{\mathrm{cut}},\Delta t_{\mathrm{step}},\Delta t_{\mathrm{ramp}})\in\Theta.$
The boundaries are
$t_{\mathrm{idle}}$ $=\Delta t_{\mathrm{idle}},$
$t_{\mathrm{peak}}$ $=t_{\mathrm{idle}}+\Delta t_{\mathrm{rise}},$
$t_{\mathrm{end}}$ $=t_{\mathrm{peak}}+\Delta t_{\mathrm{decay}},$
$t_{\mathrm{step}}$ $=t_{\mathrm{cut}}+\Delta t_{\mathrm{step}},$
$t_{\mathrm{ramp}}$ $=t_{\mathrm{step}}+\Delta t_{\mathrm{ramp}}.$
All parameters vary in bounded ranges, so $\Theta$ is compact.
Appendix B.2 Deterministic envelope
A piecewise envelope $e_{\theta}(t)$ is constructed as follows.
Idle and micro-bump.
A baseline amplitude $A_{\mathrm{idle}}$ is maintained on $0,t_{\mathrm{idle}})$, with a small pre-transient bump of the form
$b(t)=A_{\mathrm{bump}}\sin\bigl(\varphi(t)\bigr),\qquad\varphi(t)\in[0,\pi/2],$
applied over a short interval ending at $t_{\mathrm{idle}}$.
Fast rise.
On $t_{\mathrm{idle}},t_{\mathrm{peak}})$ the envelope increases linearly from the end of the bump to the peak amplitude $A_{\mathrm{peak}}$:
$e_{\theta}(t)=\text{lin}\bigl(t;\,A_{\mathrm{bump,end}},\,A_{\mathrm{peak}}\bigr).$
Shoulder and drop.
A short plateau around $A_{\mathrm{peak}}$ is followed by a linear decrease to
$A_{\mathrm{drop}}=\alpha\,A_{\mathrm{peak}},\qquad\alpha\in(0,1),$
over a compact-duration window.
Decay to plateau.
Over $t_{\mathrm{shoulder,end}},t_{\mathrm{end}})$ the envelope decays linearly to the plateau amplitude $A_{\mathrm{plateau}}$.
Fast cutoff and slow decay.
At $t_{\mathrm{cut}}$ a fast drop occurs to
$A_{\mathrm{step}}=\beta A_{\mathrm{plateau}},\qquad\beta\in(0,1),$
followed by a slow exponential decay:
$e_{\theta}(t)=A_{\mathrm{step}}e^{-k(t-t_{\mathrm{step}})},\qquad t_{\mathrm{ramp}}\leq t\leq T,$
with $k>0$ bounded.
Appendix B.3 AC carrier and observable waveform
The synthetic observable is the full AC waveform
$x_{\theta}(t)=e_{\theta}(t)\,\sin(2\pi f_{\mathrm{ac}}t+\phi)+\eta_{\theta}(t),$
where:
  • $f_{\mathrm{ac}}$ is the AC carrier frequency,
  • $\phi$ is a bounded phase,
  • $\eta_{\theta}(t)$ is a bounded perturbation term consisting of:
    $\eta_{\theta}(t)=\eta^{\mathrm{white}}_{\theta}(t)+\eta^{\mathrm{mult}}_{\theta}(t)+\eta^{\mathrm{OU}}_{\theta}(t),$
    combining small white noise, multiplicative amplitude-dependent noise, and a temporally correlated Ornstein–Uhlenbeck component.
Appendix B.4 RMS computation
No RMS envelope is generated analytically. Instead, the RMS signal used for geometric analysis is computed directly from $x_{\theta}(t)$ using the same sliding-window RMS operator applied to the real electromechanical data. This preserves complete methodological consistency between real and synthetic signals and prevents introducing artifacts from manual RMS design.
Appendix B.5 Compactness
Since $\Theta$ is compact and $\theta\mapsto x_{\theta}(\cdot)$ is continuous in the supremum norm, the synthetic manifold
$\mathcal{M}_{\mathrm{elec}}=\{x_{\theta}:\theta\in\Theta\}$
is compact in $C^{0}([0,T])$ and therefore admits a finite Hausdorff radius.
Implementation
A reference implementation of this generator, including the AC waveform, noise components, and RMS computation pipeline, will be provided in the public repository associated with this work.
Figure 11: Example RMS envelopes derived from synthetic AC traces (gray) compared with a representative real RMS signal (green). Synthetic RMS profiles are obtained by computing RMS from the AC waveform, using the exact same processing pipeline applied to the real electromechanical data. The green curve shows a random real AC RMS sample, included to illustrate the typical morphology observed in field recordings.
Appendix Appendix C Synthetic Generator for the Electrochemical Domain
Appendix C.1 Electrochemical discharge structure
Battery discharge under fixed load conditions produces voltage curves with characteristic properties:
  • monotonic or quasi-monotonic decay,
  • smooth curvature with a knee region,
  • bounded voltage range determined by chemistry and operating limits,
  • absence of high-frequency transients.
These properties reflect deterministic electrochemical constraints governed by reaction kinetics, internal resistance, and diffusion processes.
Appendix C.2 Parameterization
Each discharge curve is parameterized by a vector
$\theta=(V_{0},V_{\min},\alpha,\beta,t_{k},\gamma,\varepsilon)\in\Theta,$
where:
  • $V_{0}$ is the initial voltage,
  • $V_{\min}$ is the cutoff voltage,
  • $\alpha$ controls early-stage decay,
  • $\beta$ controls mid-stage curvature,
  • $t_{k}$ defines the knee location,
  • $\gamma$ controls terminal tapering,
  • $\varepsilon$ bounds smooth perturbations.
All parameters vary within bounded intervals, so $\Theta$ is compact.
Appendix C.3 Deterministic discharge model
The nominal discharge envelope is defined as
$v_{\theta}(t)=V_{\min}+(V_{0}-V_{\min})\exp\!\left(-\alpha t-\beta\max(0,t-t_{k})^{2}\right),\qquad t\in[0,T].$
This formulation captures:
  • an initial exponential decay,
  • a curvature increase around the knee,
  • a smooth approach to the cutoff voltage.
Appendix C.4 Smooth bounded perturbations
To account for structured variability while preserving continuity, we add a bounded smooth perturbation:
$\eta_{\theta}(t)=\varepsilon\sum_{j=1}^{J}c_{j}(\theta)\,\psi_{j}(t),$
where:
  • $\psi_{j}$ are fixed smooth basis functions on $[0,T]$,
  • $c_{j}(\theta)$ are bounded continuous coefficients,
  • $J$ is finite.
The final observable voltage curve is
$x_{\theta}(t)=v_{\theta}(t)+\eta_{\theta}(t).$
Appendix C.5 Compactness of the electrochemical manifold
The synthetic electrochemical manifold is defined as
$\mathcal{M}_{\mathrm{bat}}=\{\,x_{\theta}(\cdot):\theta\in\Theta\,\}\subset C^{0}([0,T]).$
Since $\Theta$ is compact and the map $\theta\mapsto x_{\theta}(\cdot)$ is continuous in the supremum norm, the image $\mathcal{M}_{\mathrm{bat}}$ is compact by the standard image-of-compact-is-compact argument. Consequently, the electrochemical perceptual manifold admits a finite Hausdorff radius.
Appendix C.6 Status of empirical evaluation
At the time of submission, the electrochemical domain is included as a theoretically grounded extension of the framework. Empirical saturation analysis on real battery discharge datasets is ongoing and will follow the same Monte Carlo radius estimation protocol described in Section 5.
The framework therefore makes a concrete geometric prediction: electrochemical discharge curves, despite chemical complexity, must occupy a compact perceptual manifold exhibiting early saturation of the empirical radius. This prediction will be evaluated in future work using public battery aging datasets.
Appendix Appendix D Synthetic Generators for the ECG Domain
This appendix details the synthetic generators used to construct compact functional manifolds representative of electrocardiogram (ECG) signals. The goal is not to reproduce the full electrophysiological complexity of cardiac dynamics, but to generate bounded, continuous families of waveforms with stereotyped heartbeat morphology suitable for geometric analysis.
Two generators of increasing abstraction are considered: (i) a dynamical ECG generator based on the McSharry model, and (ii) a purely morphological Gaussian generator. Both produce continuous signals on a fixed time interval and induce compact perceptual manifolds in $C^{0}([0,T])$.
Appendix D.1 Common preprocessing and representation
All synthetic ECG signals are generated on a fixed interval $[0,T]$ centered around the R-peak. Signals are resampled onto a uniform temporal grid and normalized in amplitude using the same preprocessing pipeline applied to real ECG data. This guarantees that synthetic and real signals are directly comparable under the same metric geometry.
Appendix D.2 McSharry dynamical ECG generator
The McSharry generator [10] models ECG signals using a low-dimensional nonlinear dynamical system designed to produce a stereotyped P–QRS–T morphology. While simplified, this model captures the gross temporal structure of cardiac cycles.
Appendix D.2.1 Parameterization
Each realization is determined by a parameter vector
$\theta=(A_{P},A_{Q},A_{R},A_{S},A_{T};\;t_{P},t_{Q},t_{R},t_{S},t_{T};\;\sigma_{P},\sigma_{Q},\sigma_{R},\sigma_{S},\sigma_{T};\;\varepsilon)\in\Theta,$
where:
  • $A_{k}$ control the amplitudes of the P, Q, R, S, and T components,
  • $t_{k}$ define their temporal locations,
  • $\sigma_{k}$ control their temporal widths,
  • $\varepsilon$ bounds smooth perturbations.
All parameters vary within bounded intervals, so $\Theta$ is compact.
Appendix D.2.2 Deterministic waveform construction
The nominal ECG waveform is constructed as a superposition of smooth components:
$f_{\theta}(t)=\sum_{k\in\{P,Q,R,S,T\}}A_{k}\exp\!\left(-\frac{(t-t_{k})^{2}}{2\sigma_{k}^{2}}\right).$
Additional smooth perturbations are added to account for structured variability:
$\eta_{\theta}(t)=\varepsilon\sum_{j=1}^{J}c_{j}(\theta)\,\psi_{j}(t),$
where $\psi_{j}$ are fixed smooth basis functions on $[0,T]$ and $J$ is finite.
The final observable signal is
$x_{\theta}(t)=f_{\theta}(t)+\eta_{\theta}(t).$
Appendix D.2.3 Compactness
Since $\Theta$ is compact and the map $\theta\mapsto x_{\theta}(\cdot)$ is continuous in the supremum norm, the synthetic ECG manifold generated by the McSharry model,
$\mathcal{M}_{\mathrm{ECG}}^{\mathrm{McS}}=\{x_{\theta}:\theta\in\Theta\},$
is compact in $C^{0}([0,T])$ and admits a finite Hausdorff radius.
Appendix D.3 Gaussian morphological ECG generator
To decouple geometric properties from physiological modeling assumptions, we also consider a purely morphological ECG generator based on Gaussian components. This generator ignores cardiac dynamics entirely and retains only coarse waveform structure.
Appendix D.3.1 Parameterization
Each synthetic beat is defined by
$\theta=(a_{1},a_{2},a_{3};\;t_{1},t_{2},t_{3};\;\sigma_{1},\sigma_{2},\sigma_{3};\;b,\;\varepsilon)\in\Theta,$
where:
  • $a_{k}$ control the amplitudes of three dominant excursions (e.g. P/QRS/T),
  • $t_{k}$ define their temporal locations,
  • $\sigma_{k}$ control their widths,
  • $b$ defines a baseline offset,
  • $\varepsilon$ bounds smooth perturbations.
All parameters are bounded, so $\Theta$ is compact.
Appendix D.3.2 Deterministic waveform
The deterministic component is
$f_{\theta}(t)=b+\sum_{k=1}^{3}a_{k}\exp\!\left(-\frac{(t-t_{k})^{2}}{2\sigma_{k}^{2}}\right),$
with a smooth perturbation term $\eta_{\theta}(t)$ defined as in the previous section. The observable waveform is
$x_{\theta}(t)=f_{\theta}(t)+\eta_{\theta}(t).$
Appendix D.3.3 Compactness
As before, compactness of $\Theta$ and continuity of the construction imply that the Gaussian ECG manifold
$\mathcal{M}_{\mathrm{ECG}}^{\mathrm{Gauss}}=\{x_{\theta}:\theta\in\Theta\}$
is compact in $C^{0}([0,T])$ and admits a finite Hausdorff radius.
Appendix D.4 Remarks on realism and geometry
Neither generator is intended to produce physiologically accurate ECG signals. The McSharry model captures coarse heartbeat dynamics but omits many biological details, while the Gaussian generator is purely morphological.
Nevertheless, both generators produce compact functional manifolds whose geometric saturation properties closely match those observed in real ECG data. This demonstrates that early saturation and finite perceptual radius are geometric consequences of bounded deterministic variability, not of detailed physiological realism.
Implementation
Reference implementations of both ECG generators, together with preprocessing and resampling routines identical to those applied to real ECG data, will be provided in the public repository associated with this work.